# Relative-error monotonicity testing

### v1

@@ 4.1 Warmup: A nonadaptive $O(n/\varepsilon)$-query algorithm (that is not given $N$)
{mdframed}
Algorithm 1: An edge-testing algorithm
Input: $n$, $\varepsilon$, and access to the oracles $\mathrm{MQ}(f)$ and $\mathrm{Samp}(f)$.
Output: Either “monotone” or “not monotone”.
  1. 1.
    Query the sample oracle $\mathrm{Samp}(f)$ for $a\mathrel{\mathop{\ordinarycolon}}=4n/\varepsilon$ times.
  2. 2.
    For every sample $x$ received, uniformly sample $i\in[n]$, and query the oracle $\mathrm{MQ}(f)$ on $x\oplus e_{i}$ (which is $x$ with its $i$-th bit flipped). If $x_{i}=0$ (so that $x\prec x\oplus e_{i}$) and $f(x\oplus e_{i})=0$ (which means $\{x,x\oplus e_{i}\}$ is a violating edge), then output “not monotone” and halt.
  3. 3.
    If the algorithm did not output “not monotone” in Step 2, then output “monotone”.
@@
4.2 A non-adaptive $O(\log(N)/\varepsilon)$-query algorithm that is given $N$
@@
The key to this theorem is the following observation: if $f$ is monotone and $|f^{-1}(1)|=N$, then for any $x\in f^{-1}(1)$, the value of $\|x\|_{1}$ (i.e. the Hamming weight of $x$) is at least $n-\log_{2}N$, because $f(y)=1$ for any $y\succ x$ and the number of such $y$ is $2^{n-\|x\|_{1}}\leq N$. Therefore, if we know $N$ and we found a point $x\in\{0,1\}^{n}$ with $f(x)=1$ and $\|x\|_{1}<n-\log_{2}N$, then $f$ cannot be monotone.
@@
{mdframed}
Algorithm 2: A non-adaptive algorithm that is given $N$
Input: $n$, $\varepsilon$, $N$, and access to the oracles $\mathrm{MQ}(f)$ and $\mathrm{Samp}(f)$.
Output: Either “monotone” or “not monotone”.
  1. 1.
    Query the sample oracle $\mathrm{Samp}(f)$ for $a\mathrel{\mathop{\ordinarycolon}}=16\log_{2}(N)/\varepsilon$ times.
  2. 2.
    For every sample $x$ received:
    • If $\|x\|_{1}<n-2\log_{2}N$, output “not monotone” and halt.
    • Otherwise, uniformly sample $i$ from $\{i\in[n]\mathrel{\mathop{\ordinarycolon}}x_{i}=0\}$, and query the oracle $\mathrm{MQ}(x)$ on $x\oplus e_{i}$. If $f(x\oplus e_{i})=0$ (which means $\{x,x\oplus e_{i}\}$ is a violating edge), then output “not monotone” and halt.
  3. 3.
    If the algorithm did not output “not monotone” in Step 2, then output “monotone”.
@@
Now suppose that $f$ is $\varepsilon$-relatively-far from monotone. Consider the following two cases:
  • Case 1: There are at least $\varepsilon N/2$ elements of $f^{-1}(1)$ that have Hamming weight less than $n-2\log_{2}N$. In this case, the probability that 4.2 draws a sample that has Hamming weight less than $n-2\log_{2}N$ and outputs “not monotone” is at least
    $31-\left(1-\frac{\varepsilon N/2}{N}\right)^{a}=1-\left(1-\frac{\varepsilon}{2}\right)^{16\log_{2}(N)/\varepsilon}\geq 2/$ 3 .
  • Case 2: There are at most $\varepsilon N/2$ elements of $f^{-1}(1)$ that has Hamming weight at least $n-2\log_{2}N$. In this case, if one of the $x$ sampled in Stage 1 of the algorithm has Hamming weight less than $n-2\log_{2}N$, then the algorithm outputs “not monotone”. Below we condition on the event that all $x$ sampled in the algorithm have Hamming weight at least $n-2\log_{2}N$.
@@
    $31-\left(1-\frac{\varepsilon N/4}{2N\log_{2}N}\right)^{a}=1-\left(1-\frac{\varepsilon}{8\log_{2}N}\right)^{16\log_{2}(N)/\varepsilon}\geq 2/$ 3 .
@@
4.3 An $O(\log(N)/\varepsilon)$-complexity algorithm that does not know $N$
@@
{mdframed}
Algorithm 3: An algorithm that is not given $N$
Input: $n$, $\varepsilon$, and access to the oracles $\mathrm{MQ}(f)$ and $\mathrm{Samp}(f)$.
Output: Either “monotone” or “not monotone”.
  1. 1.
    Query the sample oracle $\mathrm{Samp}(f)$ for $c\mathrel{\mathop{\ordinarycolon}}=8\log_{2}(1/\delta)/\varepsilon$ times. Let $z$ be the sample that has the least Hamming weight (if there are multiple such $z$, choose an arbitrary one); set $k\mathrel{\mathop{\ordinarycolon}}=\|z\|_{1}$. Uniformly sample $y_{1},\dots,y_{b}\in\{y\in\{0,1\}^{n}\mathrel{\mathop{\ordinarycolon}}y\succ z\}$ where $b\mathrel{\mathop{\ordinarycolon}}=4\log_{2}(1/\delta)$. If there exists $j\in[b]$ such that $f(y_{j})=0$, output “not monotone” and halt.
  2. 2.
    Query the sample oracle $\mathrm{Samp}(f)$ for $a\mathrel{\mathop{\ordinarycolon}}=16(n-k)/\varepsilon$ times.
  3. 3.
    For every sample $x$ received:
    • Uniformly sample $i$ from $\{i\in[n]\mathrel{\mathop{\ordinarycolon}}x_{i}=0\}$, and query $f(x\oplus e_{i})$. If $f(x\oplus e_{i})=0$ (which means $\{x,x\oplus e_{i}\}$ is a violating edge), output “not monotone” and halt.
  4. 4.
    If the algorithm has not output “not monotone” so far, output “monotone”.
@@
Assuming 4.3 does not output “not monotone” in Step 1 and $k$ satisfies Equation 2, we show that with probability at least $5/6$, the algorithm outputs “not monotone” in Step 3.
To this end, we bound the probability of an $x$ from $\mathrm{Samp}(f)$ leading Step 3 to return “not monotone”:
@@
    where the $\varepsilon/4$ is the probability that $x$ is in one of these $\varepsilon N/4$ violating edges and $1/(n-k)$ is the probability that the $i$ sampled in Step 3 is exactly the direction of the edge that $x$ lies in.
@@
Therefore, when $f$ is $\varepsilon$-relatively-far from monotone, 4.3 outputs “not monotone” with probability at least $2/3$.
@@
5 Warmup: A one-sided non-adaptive $\Omega(\log N)$ lower bound
As a warmup, we give a simple one-sided non-adaptive lower bound for relative-error monotonicity testing. In particular, we prove the following theorem, which is inspired by the one-sided non-adaptive monotonicity testing lower bound of [FLN+02] in the standard model:
Theorem 14.
There is a constant $\varepsilon_{0}>0$ and a parameter $N=\smash{\Theta({n\choose 3n/4})}$ such that any one-sided nonadaptive algorithm for testing whether a function $f$ with sparsity $N$ is monotone or has relative distance at least $\varepsilon_{0}$ from monotone functions must make ${\Omega}(\log N)$ many queries.
To prove this theorem, we focus on the following family of functions. For each $i\in[n]$, define $f_{i}\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow\{0,1\}$ as follows:
$f_{i}(x)=\begin{cases}1&\textrm{if }\|x\|_{1}>3n/4+1\\ 0&\textrm{if }\|x\|_{1}<3n/4\\ 1-x_{i}&\textrm{if }\|x\|_{1}\in\{3n/4,3n/4+1\}.\\ \end{cases}$
Note that all $f_{i}$ have the same sparsity $\smash{N=\Theta({n\choose 3n/4})}$.
It is easy to show that every $f_{i}$ has large relative distance from monotone functions:
Claim 15.
For every $i\in[n]$, we have $\mathsf{rel}\text{-}\mathsf{dist}(f_{i},\mathcal{C}_{\textsf{monotone}})=\Omega(1)$.
Proof.
First observe that $|f_{i}^{-1}(1)|$ can be upper bounded as
$|f_{i}^{-1}(1)|\leq\sum_{j=3n/4}^{n}\binom{n}{j}=O\bigg{(}\binom{n}{3n/4}\bigg{)}.$
Define $\mathcal{A}=\{x\mathrel{\mathop{\ordinarycolon}}x_{i}=0$ and $\|x\|_{1}=3n/4-1\}$ and $\mathcal{A}^{\uparrow}=\{x\mathrel{\mathop{\ordinarycolon}}x_{i}=1$ and $\|x\|_{1}=3n/4\}$. Observe that for each $x\in\mathcal{A}$, there is a unique $y\in\mathcal{A}^{\uparrow}$ such that $f_{i}(x)=1$ and $f_{i}(y)=0$. Further, $x$ and $y$ differ only in the $i^{th}$ coordinate. This implies that $\mathrm{dist}(f_{i},\mathcal{C}_{\textsf{monotone}})\geq|\mathcal{A}|$. Thus,
$\mathsf{rel}\text{-}\mathsf{dist}(f_{i},\mathcal{C}_{\textsf{monotone}})\geq\Theta\bigg{(}\frac{|\mathcal{A}|}{\binom{n}{3n/4}}\bigg{)}=\Theta(1),\quad\text{as was to be shown.}\qed$
Given 15 (letting $\varepsilon_{0}$ be the hidden constant), our goal is to show that any randomized, non-adaptive algorithm that can find a violation in $f_{i}$ with probability at least $2/3$, for every $i\in[n]$, must make $\Omega(n)$ queries. We start by showing that it suffices to prove the same $\Omega(n)$ lower bound for any randomized, non-adaptive algorithm that has access to the black-box oracle $\mathrm{MQ}(f)$ only.
Note that every function $f_{i}$ is a two-layer function as defined in Equation 3. Indeed all functions used in our lower bound proofs in the next two sections are two-layer functions as well.
One reason that two-layer testing is helpful for lower bound arguments is because the $\mathrm{Samp}(f)$ oracle is not needed for two-layer testing — it can be simulated, at low overhead, with the $\mathrm{MQ}$ oracle. This means that to prove a lower bound in our relative-error model for two-layer functions, it is sufficient to consider the slightly simpler setting in which the testing algorithm’s only access to the unknown two-layer function $f$ is via the $\mathrm{MQ}$ oracle.
Claim 16.
If $f\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\to\{0,1\}$ is a two-layer functions, then $q$ calls to the $\mathrm{Samp}(f)$ oracle can be simulated, with success probability at least $99/100$, by making $O(q)$ calls to the $\mathrm{MQ}$ oracle.
Proof.
The simulation works by simply making $Cq$ (for a suitable large constant $C$) many $MQ(f)$ queries on independent uniform random points $\bm{x}\in\{0,1\}^{n}$ that satisfy $\|\bm{x}\|_{1}\geq 3n/4$, and using the first $q$ points for which $f(\bm{x})=1.$ It is clear that each such point $\bm{x}$ in the sample (for which $f(\bm{x})=1$) corresponds to an independent draw from $\mathrm{Samp}(f)$; correctness follows from a simple probabilistic argument using the fact that for any two-layer function $f$, at least an $\Omega(1)$ fraction of all $n$-bit strings with Hamming weight at least $43n/$ 4 are satisfying assignments of $f$. ∎
Given 16, it suffices to show that any randomized, non-adaptive, $\mathrm{MQ}$-only algorithm that can find a violation in $f_{i}$ with probability at least $(99/100)\cdot 2/3$ must make $\Omega(n)$ $\mathrm{MQ}$ queries. To this end, letting $\mathcal{D}_{\mathrm{no}}$ be the uniform distribution on $f_{i}$, it suffices to prove Lemma 17 below, because any randomized, $q$-query, non-adaptive algorithm that can find a violation in every $f_{i}$ with probability at least $(99/100)\cdot 2/3$ implies a deterministic algorithm that finds a violation in $\bm{f}\sim\mathcal{D}_{\mathrm{no}}$ with probability at least $(99/100)\cdot 2/3$.
Lemma 17.
Any deterministic, $q$-query, non-adaptive, $\mathrm{MQ}$-only algorithm finds a violation in $\bm{f}\sim\mathcal{D}_{\mathrm{no}}$ with probability at most $q/n$.
Proof of Lemma 17.
Let $Q$ be the set of $q$ queries. We assume without loss of generality that every $x\in Q$ has $\|x\|_{1}\in\{3n/4,3n/4+1\}$. We will show that $Q$ can reveal a violation for at most $q-1$ of the functions $f_{i}$. To argue this, consider an undirected graph $G=(Q,E)$ in which the vertices correspond to queries in $Q$. Join any two vertices in $Q$ with an edge labeled by $j$ if they differ at exactly the $j^{th}$ coordinate. Next, we drop all the edges until we just retain a spanning forest $G$. Let the edges of the spanning forest be $E^{\prime}$. Note that the spanning forest has at most $q-1$ edges.
Now, consider a function $f_{i}$ whose violation is revealed by the set of queries. For this to be the case, it must mean that there is a pair of adjacent vertices — call them $u$ and $v$ – in $G$ such that the corresponding edge is labeled by $i$. This means that in the spanning forest, $u$ and $v$ must be in the same component. Furthermore, there must be at least one edge in the path connecting $u$ and $v$ in the spanning forest which is labeled by $i$. Since the number of edges in the spanning forest is at most $q-1$, it follows that $Q$ can reveal a violation for at most $q-1$ of the functions $f_{i}$. ∎
Remark 18.
In Theorem 14 as well as the rest of the paper, we work with two-layer functions as described above, where the two layers in question are $43n/$ 4 and $13n/4+$ 1. This choice is made just for concreteness to aid with readability; it is easy to verify that the definition of two-layer functions could be altered to use the two layers $\alpha n$ and $\alpha n+1$, for any constant $\alpha\in(1/2,1)$, and that Theorem 14 would still go through with $N=\Theta({n\choose\alpha n})$.
@@ 6 A two-sided non-adaptive lower bound
As it will become clear in Section 6.1, all functions used in our lower bound proof are two-layer functions. We will prove that any randomized, non-adaptive, $\mathrm{MQ}$-only algorithm for testing two-layer functions with success probability $(99/100)\cdot 2/3$ must make $\tilde{\Omega}(n)$ queries. It follows from 16 that, even with access to both $\mathrm{MQ}$ and $\mathrm{Samp}$, any randomized, non-adaptive algorithm for testing two-layer functions with success probability $2/3$ must make $\tilde{\Omega}(n)$ queries. In the rest of the section, we focus on $\mathrm{MQ}$-only algorithms.

### v2

@@ 4.1 Warmup: A nonadaptive $O(n/\varepsilon)$-query algorithm (that is not given $N$)
Algorithm 1: An edge-testing algorithm Input: $n$, $\varepsilon$, and access to the oracles $\mathrm{MQ}(f)$ and $\mathrm{Samp}(f)$. Output: Either “monotone” or “not monotone”. 1. Query the sample oracle $\mathrm{Samp}(f)$ for $a\mathrel{\mathop{\ordinarycolon}}=4n/\varepsilon$ times. 2. For every sample $x$ received, uniformly sample $i\in[n]$, and query the oracle $\mathrm{MQ}(f)$ on $x\oplus e_{i}$ (which is $x$ with its $i$-th bit flipped). If $x_{i}=0$ (so that $x\prec x\oplus e_{i}$) and $f(x\oplus e_{i})=0$ (which means $\{x,x\oplus e_{i}\}$ is a violating edge), then output “not monotone” and halt. 3. If the algorithm did not output “not monotone” in Step 2, then output “monotone”.
@@
4.2 A non-adaptive $O(\log(N)/\varepsilon)$-query algorithm that is given NN
@@
The key to this theorem is the following observation: if $f$ is monotone and $|f^{-1}(1)|=N$, then for any $x\in f^{-1}(1)$, the value of $\|x\|_{1}$ (i.e. the Hamming weight of $x$) is at least $n-\log_{2}N$, because $f(y)=1$ for any $y\succ x$. Therefore, if we know $N$ and we found a point $x\in\{0,1\}^{n}$ with $f(x)=1$ and $\|x\|_{1}<n-\log_{2}N$, then $f$ cannot be monotone.
@@
Algorithm 2: A non-adaptive algorithm that is given NN Input: $n$, $\varepsilon$, $N$, and access to the oracles $\mathrm{MQ}(f)$ and $\mathrm{Samp}(f)$. Output: Either “monotone” or “not monotone”. 1. Query the sample oracle $\mathrm{Samp}(f)$ for $a\mathrel{\mathop{\ordinarycolon}}=16\log_{2}(N)/\varepsilon$ times. 2. For every sample $x$ received: If $\|x\|_{1}<n-2\log_{2}N$, output “not monotone” and halt. Otherwise, uniformly sample $i$ from $\{i\in[n]\mathrel{\mathop{\ordinarycolon}}x_{i}=0\}$, and query the oracle $\mathrm{MQ}(x)$ on $x\oplus e_{i}$. If $f(x\oplus e_{i})=0$ (which means $\{x,x\oplus e_{i}\}$ is a violating edge), then output “not monotone” and halt. 3. If the algorithm did not output “not monotone” in Step 2, then output “monotone”.
@@
Now consider the case where $f$ is $\varepsilon$-relatively-far from monotone. Let $g$ be the closest (under Hamming distance) monotone function from $f$, and let $S=f^{-1}(1)\ \triangle\ g^{-1}(1)$, i.e. the set of points where $f$ and $g$ differ. Note that $|S|\geq\varepsilon N$. We have the following two cases:
  • Case 1: There are at least $\varepsilon N/2$ elements of $S$ that have Hamming weight less than $n-2\log_{2}N$. In this case, the probability that 4.2 draws a sample that has Hamming weight less than $n-2\log_{2}N$ and outputs “not monotone” is at least
    $1-\left(1-\frac{\varepsilon N/2}{N}\right)^{a}=1-\left(1-\frac{\varepsilon}{2}\right)^{16\log_{2}(N)/\varepsilon}\geq 2/3.$
  • Case 2: There are more than $\varepsilon N/2$ elements of $S$ that has Hamming weight at least $n-2\log_{2}N$. In this case, if one of the $x$ sampled in Stage 1 of the algorithm has Hamming weight less than $n-2\log_{2}N$, then the algorithm outputs “not monotone”. Below we condition on the event that all $x$ sampled in the algorithm have Hamming weight at least $n-2\log_{2}N$.
@@
    $1-\left(1-\frac{\varepsilon N/4}{2N\log_{2}N}\right)^{a}=1-\left(1-\frac{\varepsilon}{8\log_{2}N}\right)^{16\log_{2}(N)/\varepsilon}\geq 2/3.$
@@
4.3 An $O(\log(N)/\varepsilon)$-complexity algorithm that does not know NN
@@
Algorithm 3: An algorithm that is not given NN Input: $n$, $\varepsilon$, and access to the oracles $\mathrm{MQ}(f)$ and $\mathrm{Samp}(f)$. Output: Either “monotone” or “not monotone”. 1. Query the sample oracle $\mathrm{Samp}(f)$ for $c\mathrel{\mathop{\ordinarycolon}}=8\log_{2}(1/\delta)/\varepsilon$ times. Let $z$ be the sample that has the least Hamming weight (if there are multiple such $z$, choose an arbitrary one); set $k\mathrel{\mathop{\ordinarycolon}}=\|z\|_{1}$. Uniformly sample $y_{1},\dots,y_{b}\in\{y\in\{0,1\}^{n}\mathrel{\mathop{\ordinarycolon}}y\succ z\}$ where $b\mathrel{\mathop{\ordinarycolon}}=4\log_{2}(1/\delta)$. If there exists $j\in[b]$ such that $f(y_{j})=0$, output “not monotone” and halt. 2. Query the sample oracle $\mathrm{Samp}(f)$ for $a\mathrel{\mathop{\ordinarycolon}}=16(n-k)/\varepsilon$ times. 3. For every sample $x$ received: Uniformly sample $i$ from $\{i\in[n]\mathrel{\mathop{\ordinarycolon}}x_{i}=0\}$, and query $f(x\oplus e_{i})$. If $f(x\oplus e_{i})=0$ (which means $\{x,x\oplus e_{i}\}$ is a violating edge), output “not monotone” and halt. 4. If the algorithm has not output “not monotone” so far, output “monotone”.
@@
Assuming 4.3 does not output “not monotone” in Step 1 and $k$ satisfies Equation 2, we show that with probability at least $5/6$, the algorithm outputs “not monotone” in Step 3. To this end, we bound the probability of an $x$ from $\mathrm{Samp}(f)$ leading Step 3 to return “not monotone”:
@@
    where the $\varepsilon/4$ is the probability that $x$ is in one of these $\varepsilon N/4$ violating edges and 1/(nk)1/(n-k) is the probability that the $i$ sampled in Step 3 is exactly the direction of the edge that $x$ lies in.
@@
and therefore, when $f$ is $\varepsilon$-relatively-far from monotone, 4.3 outputs “not monotone” with probability at least $2/3$.
@@ 5 A two-sided non-adaptive lower bound
5.1 A useful class of functions for lower bounds: Two-layer functions
All functions used in our lower bound proofs in the next two sections are two-layer functions.
One reason that two-layer functions are helpful for lower bound arguments is because the $\mathrm{Samp}(f)$ oracle is not needed for two-layer testing — it can be simulated, at low overhead, with the $\mathrm{MQ}$ oracle. This means that to prove a lower bound in our relative-error model for two-layer functions, it is sufficient to consider the slightly simpler setting in which the testing algorithm’s only access to the unknown two-layer function $f$ is via the $\mathrm{MQ}$ oracle.
Claim 15.
If $f\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\to\{0,1\}$ is a two-layer functions, then for any constant $\tau>0$, making $q$ calls to the $\mathrm{Samp}(f)$ oracle can be simulated, with success probability at least $1-\tau$, by making $O(q)$ calls to the $\mathrm{MQ}$ oracle.
Proof.
The simulation works by simply making $Cq$ (for a suitable large constant $C$) many $MQ(f)$ queries on independent uniform random points $\bm{x}\in\{0,1\}^{n}$ that satisfy $\|\bm{x}\|_{1}\geq 3n/4$, and using the first $q$ points for which $f(\bm{x})=1.$ It is clear that each such point $\bm{x}$ in the sample (for which $f(\bm{x})=1$) corresponds to an independent draw from $\mathrm{Samp}(f)$; correctness follows from a simple probabilistic argument using the fact that for any two-layer function $f$, at least an $\Omega(1)$ fraction of all $n$-bit strings with Hamming weight at least $3n/4$ are satisfying assignments of $f$. ∎
Remark 16.
In Theorem 14 as well as the rest of the paper, we work with two-layer functions as described above, where the two layers in question are $3n/4$ and $3n/4+1$. This choice is made just for concreteness to aid with readability; it is easy to verify that the definition of two-layer functions could be altered to use the two layers $\alpha n$ and $\alpha n+1$, for any constant $\alpha\in(1/2,1)$, and that Theorem 14 would still go through with $N=\Theta({n\choose\alpha n})$.
@@ 5.1 A useful class of functions for lower bounds: Two-layer functions
As it will become clear in Section 5.2, all functions used in our lower bound proof are two-layer functions. We will prove that any deterministic, “two-stage” (see Definition 20) $\mathrm{MQ}$-only algorithm for testing two-layer functions with success probability $(99/100)\cdot 2/3$ must make $\tilde{\Omega}(n)$ queries. As discussed at the end of Section 5.2, this implies Theorem 14.